Find the general solution of each differential equation, expressing y explicitly in terms of x.

3) dy/dx= 3x2(y+1)

4) dy/dx = y-1 / 2x-1




1
Expert's answer
2021-09-14T06:09:43-0400

(i)

"\frac{dy}{dx} = 3x²(y+1)"

=> "\frac{dy}{y+1} = 3x²dx"

Integrating we get

"\int\frac{dy}{y+1} = \int3x²dx"

=> ln |y+1| = x³ + ln |A| where ln |A| is integration constant

=> ln |y+1| - ln |A| = x³

=> "ln \mid {\frac {y+1}{A}}\mid = x³"

=> y + 1 = Ae

=> y = Ae - 1

This is the solution of the given differential equation

(ii)

"\frac{dy}{dx} = \frac{y-1}{2x-1}"

=> "\frac{dy}{y-1} = \frac{dx}{2x-1}"

Integrating we get

=> "\int\frac{dy}{y-1} =\frac{1}{2}\int \frac{2dx}{2x-1}"

=> ln |y-1| = "\frac{1}{2} ln |2x-1| + ln |A|"

where ln |A| is integration constant

=> ln |y-1| - ln |A| = "\frac{1}{2} ln|2x-1|"

=> "ln \mid\frac{y-1}{A}\mid= ln{\sqrt{2x-1}}"

=> "\frac{y-1}{A}= \sqrt{2x-1}"

=> y - 1 = A"\sqrt{2x-1}"

=> y = 1 + A"\sqrt{2x-1}"

This is the solution of the given differential equation